OpenStudy (anonymous):

Let I=∫ln x dx since there is no function before ln x so we can write 1 before ln x =∫lnx dx -∫{d/dx (1) ∫ln x dx} dx =1/x-∫0.1/x dx =1/x I=1/x is this right?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

SORRy the answer is 1/x +c

OpenStudy (anonymous):

@nincompoop nino is this right??

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zepdrix Help :(

zepdrix (zepdrix):

No no no. You did your parts backwards. D:

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits (1)\ln x\;dx\]U=1 V=ln x

zepdrix (zepdrix):

Woops*

zepdrix (zepdrix):

U=ln x V=1

OpenStudy (anonymous):

thats what i told to my teacher but he said before ln x 1 must be placed

zepdrix (zepdrix):

Yes I placed a before the ln x... right? It doesn't matter which order they are written, that doesn't dictate which one is U and the other is V.

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits U Vdx\quad=\quad \int\limits VUdx\]It doesn't matter if you write the 1 before or after. If I'm understanding you correctly.

zepdrix (zepdrix):

You just have to choose the correct parts. U=ln x V=1 for this problem.

OpenStudy (anonymous):

thank you very much @zepdrix :)